solve for x and simplify:
4x ln (1/4)-xln2= -6(x+1)ln(1/8)

Answer by ikleyn(53763)

.
solve for x and simplify:
4x ln(1/4)-xln2= -6(x+1)ln(1/8)
----------------------------------------

4x ln(1/4)-xln2= -6(x+1)ln(1/8).


Do you know that ln(1/4) = -2*ln(2), ln(1/8) = -3*ln(2) ?


Substitute it into your equation and then cancel the factor ln(2) in both side.


Then you will be on an easy track.